OpenStudy (anonymous):

Can some help me with a math problem i am stuck on? Here is the problem A local fish and game commission releases 500 brown trout in a new man-made lake. The birth rate is expected to increase population approximately 55 fish per year. The death rate is expected to decrease the population approximately 35 fish per year. Use a function P(t) where t is the number of years to find the population after 5 years. The population after 5 years is about ---- fishes

hero (hero):

P(t) = 500+(55-35)t = 500 + 20t

hero (hero):

P(5) = 500 + 20(5) P(5) = 500 + 100 P(5) = 600

OpenStudy (anonymous):

oh that how you do that

OpenStudy (anonymous):

I that i did something wrong

hero (hero):

It was magic

OpenStudy (anonymous):

But when i did it it said it was wrong

hero (hero):

The answer is wrong?

OpenStudy (anonymous):

It came up with 2600 and it said it was wrong

OpenStudy (anonymous):

because after you do all the first work and get your answer you have to multiply by 5 am i right

hero (hero):

I used 5 for t

OpenStudy (anonymous):

what did u get

hero (hero):

I did the work above. I'm verifying it. Hold on a sec

hero (hero):

The answer is definitely 600

hero (hero):

There's no other way to approach this

hero (hero):

You must be inputting something wrong when you're typing in the answer

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

i got another math problem You join a DVD rental club. Each time you visit the store, you can rent one DVD for $6 and additional DVDs for $4.50 each. Letting x represent the number of additional DVDs that you rent, write a function R(x) that gives the cost of renting DVDs. Then use the function to find the cost if you rent three additional DVDs.

hero (hero):

Are you taking an online test or something?

hero (hero):

If you are, I would like to help you

OpenStudy (anonymous):

sorta in an online math help things... i just trying to do so manything at one and it is hard

hero (hero):

I honestly think you're inputting a correct answer wrongly. If I were to see how you were inputting your answers, I may be able to help. Can you upload a screenshot of your previous answer and how you entered it?

OpenStudy (anonymous):

not at the moment but can you look at my other problem